# Representation Learning with Large Language Models for Recommendation

### v3

@@ Abstract.
Recommender systems have seen significant advancements with the influence of deep learning and graph neural networks, particularly in capturing complex user-item relationships. However, these graph-based recommenders heavily depend on ID-based data, potentially disregarding valuable textual information associated with users and items, resulting in less informative learned representations. Moreover, the utilization of implicit feedback data introduces potential noise and bias, posing challenges for the effectiveness of user preference learning. While the integration of large language models (LLMs) into traditional ID-based recommenders has gained attention, challenges such as scalability issues, limitations in text-only reliance, and prompt input constraints need to be addressed for effective implementation in practical recommender systems. To address these challenges, we propose a model-agnostic framework RLMRec that aims to enhance existing recommenders with LLM-empowered representation learning. It proposes a recommendation paradigm that integrates representation learning with LLMs to capture intricate semantic aspects of user behaviors and preferences. RLMRec incorporates auxiliary textual signals, develops a user/item profiling paradigm empowered by LLMs, and aligns the semantic space of LLMs with the representation space of collaborative relational signals through a cross-view alignment framework. This work further establish a theoretical foundation demonstrating that incorporating textual signals through mutual information maximization enhances the quality of representations. In our evaluation, we integrate RLMRec with state-of-the-art recommender models, while also analyzing its efficiency and robustness to noise data. Our implementation codes are available at https://github.com/HKUDS/RLMRec.
*** Corresponding author, chaohuang75@gmail.com
copyright: none
@@ 1. Introduction
However, it is important to acknowledge that recent graph-based recommenders heavily rely on ID-corresponding information for learning. In this line, the training data consists exclusively of mapped user/item indices, and their interactions are represented in an interaction matrix using binary values (1 indicating an interaction and 0 indicating no interaction). While this data arrangement has demonstrated effectiveness, one significant limitation is that it primarily relies on ID-based information, thereby potentially overlooking other valuable data, such as rich textual information associated with users and items. The absence of this additional information can lead to reduced informativeness in the learned representations. Furthermore, it is worth noting that a substantial portion of the data in these graph-based recommenders consists of implicit feedback (Rendle et al., 2012; Wang et al., 2021a), which can introduce noise from false negatives or bias (e.g., misclicks (Wang et al., 2021b) or popularity bias (Chen et al., 2023)). Consequently, the learned representations of these GNN-based models heavily rely on the inherent quality of the data. This heavy reliance on the data quality poses a potential challenge as it can lead to detrimental representations that hinder the effectiveness of recommendation systems, especially when the data contains noise.
@@
In recent times, there have been several endeavors to leverage diverse data modalities in order to enhance traditional ID-based recommenders (Yuan et al., 2023; Li et al., 2023; Geng et al., 2022). Particularly interesting is the emergence of large language models (LLMs) like GPT-4 (OpenAI, 2023) and LLaMA (Touvron et al., 2023), which have demonstrated impressive capabilities in neural language understanding tasks. This development has sparked significant interest among researchers, who are actively exploring how LLMs, with their proficiency in handling textual content, can extend the capabilities of recommendation systems beyond the original data (Fan et al., 2023; Lin et al., 2023; Liu et al., 2023b). A primary focus of current research in this field revolves around aligning recommendation approaches with the characteristics of language models through prompt design. Methods like InstructRec (Zhang et al., 2023) structure the recommendation task in an instruction-question-answering format, enabling LLMs to simultaneously address the recommendation objective and respond to intricately designed questions (Geng et al., 2022; Bao et al., 2023). However, despite displaying some recommendation capabilities, these methods still fall behind existing recommenders in terms of efficiency and precision. This can be attributed to inherent shortcomings associated with this approach, including the following key aspects:
@@
i) Scalability issues in practical recommenders. The utilization of large language models (LLMs) in personalized user behavior modeling, comes with inherent computational demands. As the scale of user behavior data increases, the computational requirements and associated inference time costs also tend to rise significantly. For instance, in the case of TALLRec (Bao et al., 2023), where recommendations are generated based on an instruction-question-answering format, the response time for LLaMA2-13B to provide recommendations to individual users stands at approximately 3.6 seconds, based on an input size of around 800 tokens (equivalent to approximately 5 users). However, this poses significant challenges when attempting to scale up the LLM-based approach for practical recommender systems with a substantial user base and extensive item catalog.
@@
To validate the aforementioned limitations, we conduct an evaluation to assess the effectiveness of directly using LLM in enhancing the re-ranking task (Hou et al., 2023; Sun et al., 2023) for recommendation on the Amazon dataset. Specifically, we utilize LightGCN (He et al., 2020) as the underlying backbone recommender model, which generate a ranking list of 50 candidate items preferred by each user based on learned user-item interaction probabilities. To further refine the recommendations, we integrate the textual information of each item using our custom prompts (for prompt format details, please refer to Appendix A.3). These prompts are then processed by ChatGPT platform (i.e., gpt-3.5-turbo). The objective of this evaluation is to identify the Top-10 and Top-20 most relevant items from the ranking list generated by LightGCN for each user through the re-ranking task.
@@
It is evident from the results in Figure 1 that the recommendations refined by the ChatGPT perform worse than the original results provided by LightGCN. This indicates that there are limitations when blindly using LLMs to improve the re-ranking process in recommendation systems that heavily rely on textual information. These limitations can be attributed to three key factors: i) The hallucination issue of LLMs, where they may suggest recommended items that are not included in the candidate set; ii) The lack of a comprehensive global text-based collaborative relationship input due to the token limit of LLMs; iii) Additionally, it is worth noting that the reranking process using LLM takes several hours to complete, which poses a challenge when dealing with large-scale data in real-world recommendation scenarios. Due to page limit, we delve deeper into this experiment and present real cases to illustrate the hallucination phenomenon for the reranking in the Appendix.
@@
Contributions. In light of the aforementioned limitations, our work aims to leverage the power of LLMs to seamlessly enhance existing recommender systems. To accomplish this, we propose a model-agnostic framework called RLMRec (Representation Learning with Large Language Models for Recommendation). The core idea of RLMRec is to utilize representation learning as a bridge between ID-based recommenders and LLMs. Our new recommendation paradigm aims to preserve the accuracy and efficiency of existing recommenders while harnessing the powerful text comprehension capabilities of LLMs to understand the intricate semantic aspects of user behaviors and preferences. To begin, we lay the theoretical groundwork by modeling the benefits of incorporating auxiliary textual signals for representation learning. This involves transforming the textual signals into meaningful representations and establishing a theoretical foundation for maximizing mutual information within general recommendation models. Furthermore, we develop a user/item profiling paradigm empowered by LLMs. This paradigm enables us to incorporate the comprehensive semantic understanding derived from LLMs into the encoded representations of users and items. By leveraging the global knowledge space of LLMs, we enhance the representation expressiveness.
@@
Furthermore, we propose to align the semantic space of LLMs and the representation space of collaborative relational signals through a cross-view alignment framework. This alignment is achieved through a cross-view mutual information maximization scheme, which allows us to find a common semantic subspace where the textual and collaborative relational embeddings are well aligned from the contrastive and generative modeling, respectively. This typically involves jointly optimizing the recommendation model parameters and the cross-modal alignment paradigm. In a nutshell, our main contributions can be summarized as follows:
@@
  • We establish a theoretical foundation to demonstrate the effectiveness of incorporating textual signals in enhancing the representation learning of existing recommenders. By utilizing mutual information maximization as the optimization direction, we show how textual signals can improve the representation quality.
@@ 2. Related Work
GNN-enhanced Collaborative Filtering. Collaborative Filtering (CF), which is a fundamental technique in recommendation systems, has been extensively studied over the years (Su and Khoshgoftaar, 2009; Koren et al., 2021). Recently, an emerging research direction involves leveraging historical user-item interactions to construct a bipartite graph and utilizing graph neural networks (GNNs) to capture high-order collaborative relationships. These graph-based methods, such as NGCF (Wang et al., 2019), GCCF (Chen et al., 2020), LightGCN (He et al., 2020), have demonstrated state-of-the-art performance, improving recommendation effectiveness. However, the sparsity and noise in implicit feedback data pose challenges to graph-based methods. To address these challenges, researchers have started exploring the use of self-supervised learning (SSL) techniques as auxiliary learning objectives to enhance robustness in recommendations (Yu et al., 2023; Yang et al., 2023). Among various SSL techniques, contrastive learning has emerged as a prominent solution in collaborative filtering models. Methods like SGL (Wu et al., 2021), SimGCL (Yu et al., 2022), NCL (Lin et al., 2022), LightGCL (Cai et al., 2023) leverage contrastive data augmentation to improve recommendation performance. In this work, we take a step further by integrating LLMs with existing CF models to effectively align the knowledge and reasoning abilities of LLMs with the collaborative relation learning for enhancing recommendation performance.
Large Language Models for Recommendation. Recently, there has been a growing interest in leveraging LLMs for recommendation systems. (Fan et al., 2023; Lin et al., 2023; Liu et al., 2023b; Wu et al., 2023). Several studies have leveraged LLMs as inference models by designing prompts that align them with recommendation tasks. For example, P5 (Geng et al., 2022) converts the user interaction data into textual prompts using item indexes, which are then used for language model training. Chat-REC (Gao et al., 2023) builds a conversational recommender by transforming user profiles and interactions into prompts for LLMs to generate recommendations. InstructRec (Zhang et al., 2023) and TALLRec (Bao et al., 2023) employ instructional designs to define recommendation tasks and fine-tune LLMs to align with these instructions for generating recommendations. However, using LLMs directly as inference models for recommendation tasks presents challenges, such as high computational costs and slow inference times. These challenges hinder the practical deployment of such models in real-world recommender systems. To address this gap, the proposed approach adopts a theoretically grounded paradigm of mutual information maximization to align the knowledge of LLMs with collaborative relation modeling, enabling scalable and effective recommendations.
@@ 4.1.3. Base Models
We evaluate the effectiveness of our RLMRec by integrating it with state-of-the-art representation-based recommenders based on SSLRec (Ren et al., 2023a). This approach allows us to assess its model-agnostic performance gain in comparison to base models.
@@ 4.1.4. Implementation Details
To ensure a fair comparison, all baseline models are evaluated using the same evaluation protocol. The dimension of representations (i.e., x and e) is set to 32 for all base models. We determine the hyperparameters specific to each model through grid search under various settings. To generate user and item profiles, we leverage the ChatGPT model (specifically, gpt-3.5-turbo) provided by OpenAI. We use the text-embedding-ada-002 (Neelakantan et al., 2022) to generate semantic representations s. During training, all methods are trained with a fixed batch size of 4096 and a learning rate of 1e-3 using the Adam optimizer. We adopt the early stop technique based on the model’s performance on the validation set. For detailed implementation of our RLMRec, we provide the source code for reference.
@@ 4.2. Performance Comparison (RQ1)
  • Overall, we consistently observe that integrating RLMRec with the backbone recommenders leads to improved performance compared to the original versions. This provides compelling evidence for the effectiveness of RLMRec. We attribute these improvements to two key factors: i) RLMRec enables the accurate user/item profiling paradigm empowered by LLMs, enhancing the representation of rich semantic information derived from user interaction behaviors. ii) Through our cross-view mutual information maximization, the CF-side relational embeddings and LLM-side semantic representations work cooperatively to enhance each other. This collaborative effort effectively filters out irrelevant noise in the recommendation features.
@@ 4.3. Ablation Study (RQ2)
  • After randomly rearranging the semantic representations to disrupt the correlation between collaborative and semantic signals, we observe a decrease in performance for both RLMRec-Con and RLMRec-Gen on the evaluated backbone models compared to their original performance. This suggests that the shuffled representations introduce noise due to the mismatch between semantic and collaborative information. It provides evidence that accurate alignment between the semantic knowledge of the LLM and collaborative relationships among users is crucial for enhancing recommendation performance.
@@ 4.4.2. Performance in Pre-training Scenarios (RQ4)
We investigate the potential of our semantically involved training mechanism as a pre-training technique to provide informative representations as initial embeddings for downstream models. Using the Yelp dataset, we utilize data from 2012 to 2017 for pre-training and divide the data from 2018 to 2019 into a training set, a validation set, and a test set (the downstream dataset). Both datasets contain the same users and items. We train vanilla LightGCN and our model on the pre-training dataset. The learned parameters are used to initialize the embeddings for vanilla LightGCN, which is then trained on the downstream dataset. Key findings from Table 3 are:
@@ 4.5. Case Study
In our investigation, we explore the integration of LLM-enhanced semantics to capture global user relationships that are not easily captured through direct message passing on the graph. Figure 6 presents a case study where the distance between user $u_{1998}$ and $u_{227}$ exceeds 3 hops. To evaluate the models’ ability to capture their relationship, we examine the similarity of user representations. We compared LightGCN and RLMRec-Con, both using the same backbone. Two metrics were introduced: a relevance score for user $u_{1998}$ and the ranking of its long-distance neighbors (greater than 3 hops) based on the score. By incorporating semantic information, derived from language models that highlight shared interests between $u_{1998}$ and $u_{227}$ (e.g., their preference for friendly service), both the relevance score and ranking increased. This indicates that the learned representations from RLMRec effectively capture global collaborative relationships beyond ID-based recommendation techniques.

### v5

@@ Abstract.
Recommender systems have seen significant advancements with the influence of deep learning and graph neural networks, particularly in capturing complex user-item relationships. However, these graph-based recommenders heavily depend on ID-based data, potentially disregarding valuable textual information associated with users and items, resulting in less informative learned representations. Moreover, the utilization of implicit feedback data introduces potential noise and bias, posing challenges for the effectiveness of user preference learning. While the integration of large language models (LLMs) into traditional ID-based recommenders has gained attention, challenges such as scalability issues, limitations in text-only reliance, and prompt input constraints need to be addressed for effective implementation in practical recommender systems. To address these challenges, we propose a model-agnostic framework RLMRec that aims to enhance existing recommenders with LLM-empowered representation learning. It proposes a recommendation paradigm that integrates representation learning with LLMs to capture intricate semantic aspects of user behaviors and preferences. RLMRec incorporates auxiliary textual signals, employs LLMs for user/item profiling, and aligns the semantic space of LLMs with collaborative relational signals through cross-view alignment. This work further demonstrates the theoretical foundation of incorporating textual signals through mutual information maximization, which improves the quality of representations. Our evaluation integrates RLMRec with state-of-the-art recommender models, while also analyzing its efficiency and robustness to noise data. Implementation codes are available at https://github.com/HKUDS/RLMRec.
Large Language Models, Recommendation, Alignment
journalyear: 2024copyright: acmlicensedconference: Proceedings of the ACM Web Conference 2024; May 13–17, 2024; Singapore, Singaporebooktitle: Proceedings of the ACM Web Conference 2024 (WWW ’24), May 13–17, 2024, Singapore, Singaporedoi: 10.1145/3589334.3645458isbn: 979-8-4007-0171-9/24/05ccs: Information systems Recommender systems
@@ 1. Introduction
However, it’s important to note that recent graph-based recommenders heavily rely on ID-corresponding information for learning. The training data in this line only consists of mapped user/item indices, and their interactions are represented in an interaction matrix using binary values (1 for interaction and 0 for no interaction). While this data arrangement has shown effectiveness, one limitation is its primary reliance on ID-based information, potentially overlooking other valuable data such as rich textual information associated with users and items. The absence of this additional information can lead to reduced informativeness in the learned representations. Furthermore, it is worth noting that a substantial portion of the data in these graph-based recommenders consists of implicit feedback (Rendle et al., 2012; Wang et al., 2021a), which can introduce noise from false negatives or bias (e.g., misclicks (Wang et al., 2021b) or popularity bias (Chen et al., 2023)). Consequently, the learned representations of these GNN-based models heavily rely on the inherent quality of the data. This heavy reliance on the data quality poses a potential challenge as it can lead to detrimental representations that hinder the effectiveness of recommendation systems, especially when the data contains noise.
@@
In recent times, there have been several endeavors to leverage diverse data modalities in order to enhance traditional ID-based recommenders (Yuan et al., 2023; Li et al., 2023; Geng et al., 2022). Particularly interesting is the emergence of large language models (LLMs) like GPT-4 (OpenAI, 2023) and LLaMA (Touvron et al., 2023), which have demonstrated impressive capabilities in neural language understanding tasks. This development has sparked significant interest among researchers, who are actively exploring how LLMs, with their proficiency in handling textual content, can extend the capabilities of recommendation systems beyond the original data (Fan et al., 2023; Lin et al., 2023; Liu et al., 2023b). A primary focus of current research in this field revolves around aligning recommendation approaches with the characteristics of language models through prompt design. Methods like InstructRec (Zhang et al., 2023) structure the recommendation task in an instruction-question-answering format, enabling LLMs to simultaneously address the recommendation objective and respond to intricately designed questions (Geng et al., 2022; Bao et al., 2023). However, these methods still lag behind existing recommenders in terms of efficiency and precision. This can be attributed to inherent shortcomings associated with this approach, including the following key aspects:
@@
i) Scalability issues in practical recommenders. Utilizing large language models (LLMs) in personalized user behavior modeling requires significant computational resources. As the scale of user behavior data grows, so do the computational demands and associated inference time costs. For instance, in TALLRec (Bao et al., 2023), where recommendations are generated based on an instruction-question-answering format, the response time for LLaMA2-13B to provide recommendations to individual users is approximately 3.6 seconds, with an input size of around 800 tokens (equivalent to approximately 5 users). However, this poses significant challenges when attempting to scale up the approach for practical recommender systems with a substantial user base and extensive item catalog.
@@
To validate the aforementioned limitations, we evaluate the effectiveness of directly using LLMs in enhancing the re-ranking task (Hou et al., 2023; Sun et al., 2023) for recommendation on the Amazon dataset. Specifically, we utilize LightGCN (He et al., 2020) as the backbone model, which retrieve a ranking list of 50 candidate items for each user. To further refine the recommendations, we integrate the textual information of each item with our custom prompts (for details, please refer to Appendix A.3). These prompts are then processed by ChatGPT (i.e., gpt-3.5-turbo). The objective of this task is to re-rank the item list for each user and identify the Top-10&20 most relevant items.
@@
It is evident from the results in Figure 1 that the recommendations refined by the ChatGPT perform worse than the original results provided by LightGCN. This indicates limitations when blindly using LLMs to improve the re-ranking process in recommendation. These limitations can be attributed to three factors: i) The hallucination issue of LLMs, suggesting items not in the candidate set; ii) The lack of a comprehensive global text-based collaborative relationship input due to token limits; iii) Additionally, it is worth noting that the reranking process using LLM takes several hours to complete, which poses a challenge when dealing with large-scale data in real-world recommendation scenarios. Due to page constraints, we provide detailed analysis and examples in the Appendix to demonstrate the phenomenon of hallucination.
@@
Contributions. In light of the aforementioned limitations, we aim to leverage the power of LLMs to seamlessly enhance existing recommender systems. To accomplish this, we propose a model-agnostic framework called RLMRec (Representation Learning with Large Language Models for Recommendation). The core idea of RLMRec is to utilize representation learning as a bridge between ID-based recommenders and LLMs. Our new recommendation paradigm aims to preserve the accuracy and efficiency of existing recommenders while harnessing the powerful text comprehension capabilities of LLMs to understand the intricate semantic aspects of user behaviors and preferences. To begin, we lay the theoretical groundwork by modeling the benefits of incorporating auxiliary textual signals for representation learning. This involves transforming the textual signals into meaningful representations and establishing a theoretical foundation for maximizing mutual information within general recommendation models. Moreover, we develop a user/item profiling paradigm empowered by LLMs, enhancing representation expressiveness by incorporating comprehensive semantic understanding from the global knowledge space of LLMs.
@@
Furthermore, we propose to align the semantic space of LLMs and the representation space of collaborative relational signals through a cross-view alignment framework. This alignment is achieved through a cross-view mutual information maximization scheme, which allows us to find a common semantic subspace where the textual and collaborative relational embeddings are well aligned from the contrastive and generative modeling, respectively. In a nutshell, our main contributions can be summarized as follows:
@@
  • We establish a theoretical foundation to demonstrate the effectiveness of incorporating textual signals in enhancing the representation learning. By utilizing mutual information maximization, we show how textual signals can improve the representation quality.
@@ 2. Related Work
GNN-enhanced Collaborative Filtering. Collaborative Filtering (CF), a fundamental technique in recommendation systems, has been extensively studied (Su and Khoshgoftaar, 2009; Koren et al., 2021). An emerging direction is to use historical user-item interactions to create a bipartite graph and employ Graph Neural Networks (GNNs) to capture high-order collaborative relationships. These graph-based methods, such as NGCF (Wang et al., 2019), GCCF (Chen et al., 2020), LightGCN (He et al., 2020), have demonstrated state-of-the-art performance, improving recommendation effectiveness. However, the sparsity and noise in implicit feedback data pose challenges to graph-based methods. To address this, researchers have explored the use of self-supervised learning (SSL) techniques as auxiliary learning objectives to enhance robustness in recommendations (Yu et al., 2023; Yang et al., 2023). Among various SSL techniques, contrastive learning has emerged as a prominent solution in collaborative filtering models. Methods like SGL (Wu et al., 2021), SimGCL (Yu et al., 2022), NCL (Lin et al., 2022), LightGCL (Cai et al., 2023) leverage contrastive data augmentation to improve recommendation performance. In this work, we take a step further by integrating LLMs with existing CF models to effectively align the knowledge and reasoning abilities of LLMs with the collaborative relation learning for enhancing recommendation performance.
Large Language Models for Recommendation. leveraging LLMs for recommendation systems has gained interest (Fan et al., 2023; Lin et al., 2023; Liu et al., 2023b; Wu et al., 2023). Several studies have leveraged LLMs as inference models by designing prompts that align them with recommendation tasks. For example, P5 (Geng et al., 2022) converts the user interaction data into textual prompts using item indexes, which are then used for language model training. Chat-REC (Gao et al., 2023) builds a conversational recommender by transforming user profiles and interactions into prompts for LLMs to generate recommendations. InstructRec (Zhang et al., 2023) and TALLRec (Bao et al., 2023) employ instructional designs to define recommendation tasks and fine-tune LLMs to align with these instructions for generating recommendations. However, using LLMs directly for recommendation tasks faces challenges like high computational costs and slow inference times. To address this, our approach adopts mutual information maximization to align LLMs knowledge with collaborative relation modeling, enabling scalable and effective recommendations.
@@ 4.1.3. Base Models
We evaluate the effectiveness of our RLMRec by integrating it with state-of-the-art representation-based recommenders based on SSLRec (Ren et al., 2023a).
@@ 4.1.4. Implementation Details
The dimension of representations (i.e., x and e) is set to 32 for all base models. We determine the hyperparameters for each model through grid search. To generate user and item profiles, we leverage the ChatGPT model (specifically, gpt-3.5-turbo) provided by OpenAI. We use the text-embedding-ada-002 (Neelakantan et al., 2022) to generate semantic representations s. During training, all methods are trained with a fixed batch size of 4096 and a learning rate of 1e-3 using the Adam optimizer. We adopt the early stop technique based on the model’s performance on the validation set.
@@ 4.2. Performance Comparison (RQ1)
  • Overall, we consistently observe that integrating RLMRec with the backbone recommenders leads to improved performance compared to the original versions. This provides compelling evidence for the effectiveness of RLMRec. We attribute these improvements to two key factors: i) RLMRec enables accurate user/item profiling empowered by LLMs, enhancing the representation of rich semantic information from user interaction behaviors. ii) our cross-view mutual information maximization facilitates the cooperative enhancement of CF-side relational embeddings and LLM-side semantic representations, effectively filtering out irrelevant noise in the recommendation features.
@@ 4.3. Ablation Study (RQ2)
  • After randomly rearranging the semantic representations to disrupt the correlation between collaborative and semantic signals, we observe a decrease in performance for both RLMRec-Con and RLMRec-Gen on the evaluated backbone models. This indicates that the shuffled representations introduce noise due to the mismatch between semantic and collaborative information. It provides evidence that accurate alignment between the LLM’s semantic knowledge and collaborative relationships among users is crucial for enhancing recommendation performance.
@@ 4.4.2. Performance in Pre-training Scenarios (RQ4)
We explore the potential of our semantically involved training mechanism as a pre-training technique for downstream models. Using the Yelp dataset, we utilize data from 2012 to 2017 for pre-training and divide the data from 2018 to 2019 into a training set, a validation set, and a test set (the downstream dataset). Both datasets contain the same users and items. We train vanilla LightGCN and our model on the pre-training dataset. The learned parameters are used to initialize the embeddings for vanilla LightGCN, which is then trained on the downstream dataset. Key findings from Table 3 are:
@@ 4.5. Case Study
We explore the integration of LLM-enhanced semantics to capture global user relationships that are not easily captured through direct message passing. Figure 6 presents a case study where the distance between user $u_{1998}$ and $u_{227}$ exceeds 3 hops. To evaluate the models’ ability to capture their relationship, we examine the similarity of user representations. We compared LightGCN and RLMRec-Con, both using the same backbone. Two metrics were introduced: a relevance score for user $u_{1998}$ and the ranking of its long-distance neighbors ($>$ 3 hops) based on the score. By incorporating semantic information derived from LLMs, such as shared interests between $u_{1998}$ and $u_{227}$ (e.g., friendly service), we observed an increase in both the relevance score and ranking. This suggests that the learned representations from RLMRec capture global collaborative relationships beyond ID-based recommendation techniques.